OpenStudy (anonymous):

WILL FAN AND MEDAL IF YOU HELP ME What is the solution to 2.3a = 9.2? A. a = 21.16 B. a = 11.5 C. a = 6.9 D. a = 4

OpenStudy (pinklion23):

ill help it is easy

OpenStudy (anonymous):

ok

OpenStudy (pinklion23):

divide 9.2 by 2.3

OpenStudy (anonymous):

oh thx and thxs for not giving me the deract answer

OpenStudy (pinklion23):

yw and thx for being my fan

OpenStudy (anonymous):

can you help with more

OpenStudy (pinklion23):

of couse

OpenStudy (anonymous):

The first one would be d. right

OpenStudy (pinklion23):

yes

OpenStudy (anonymous):

What is the solution to 19 = z + 12? z =______

OpenStudy (pinklion23):

19-12=z

OpenStudy (pinklion23):

then check your work by adding z with 12

OpenStudy (anonymous):

So easy why am I asking for help on this one

OpenStudy (pinklion23):

it is okay im hear im not leaving lol

OpenStudy (pinklion23):

so what did you get

OpenStudy (anonymous):

I am about to sumit my test lets see what I get

OpenStudy (pinklion23):

okay im ready

OpenStudy (anonymous):

100% thank you sooo much

OpenStudy (pinklion23):

and what grade you in and yw glad to help

OpenStudy (anonymous):

Your awsome

OpenStudy (pinklion23):

thx but you will get better to silly your smart score is 52

OpenStudy (anonymous):

do you do home school

OpenStudy (pinklion23):

yes

OpenStudy (anonymous):

and LOL

OpenStudy (anonymous):

oh which one I do k12

OpenStudy (pinklion23):

SAME! im in 7th grade you?

OpenStudy (anonymous):

6th grade LOL

OpenStudy (pinklion23):

hey that is awesome well i have to leave but ill be back just tag me in any other questions okay well bye take care

OpenStudy (anonymous):

you to bye